Calculate the Molarity of each of the following solutions:



d.185 g of sucrose, C12H22O11 in 1.00 L of solution



e.6.30 g of HNO3 dissolved in 255 ml of solution.



Please put what letter you answer in every solution



D: solution


E: solution

1
Expert's answer
2022-03-10T07:49:03-0500

D) Molar mass of sucrose= 342.3g/mol

moles = 185/342.3 = 0.54 mol

Molarity = moles of solute ÷ litres of solution.

Molarity = 0.54mol ÷ 1L

= 0.54M


E) Molar mass of HNO3 = 63.01g/mol

moles = 6.30 ÷ 63.01 = 0.1 mol

Molarity= moles of solute÷litres of solution

Molarity= 0.1mol ÷ 0.255L

= 0.39M